Two charged particles apply an electric force of 5.2×10 −3 N on each other. Distance between them gets twice as much as before. What will be Coulomb's force?


1
Expert's answer
2021-06-11T11:32:56-0400

Gives

"F_1=5.2\times10^{-3} N"

"F_2=?"

"r_2=2r_1"

"F_1=\frac{kq_1q_2}{r_1^2}=5.2\times10^{-3}N"

"r_2=2r_1"

"\frac{F_2}{F_1}=\frac{\frac{kq_1q_2}{r_2^2}}{\frac{kq_1q_2}{r_1^2}}"

"\frac{F_2}{F_1}=\frac{r_1^2}{(2r_2)^2}=\frac{1}{4}"

"F_2=\frac{F_1}{4}=\frac{5.2\times 10^{-3}}{4}"

"F_2=1.3\times10^{-3}N"